OpenStudy (anonymous):

Okay it will not go insane in here. A jar contains 3 red, 4 blue, and 5 white marbles. What is the probability of drawing first a red and then, without replacement, a white marble? 5/44 5/48 31/44

OpenStudy (anonymous):

\[Probability = \frac{\text{Number of Red Marbles}}{\text{Total Number of Marbles}} \times \frac{\text{Number of White Marbles}}{\text{Total Number of Marbles} - 1}\]

OpenStudy (amistre64):

also, hypergeometric would work

OpenStudy (anonymous):

3/12 * 5/12?

OpenStudy (anonymous):

3/12x5/11 = 5/44

OpenStudy (anonymous):

is it a proportion or multiply across?

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

Remember that it's the total number of marbles - 1

OpenStudy (anonymous):

For the second fraction.

OpenStudy (amistre64):

hyper might not work ... youve got more than 2 sets, i gotta read better

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

(3/12)x(5/11)= 5/44

OpenStudy (anonymous):

now you both have medals :)